After spending 1/4 of my money, then 1/5 of the remainder, I had $66 left. How many dollars did I start with?
Answer by htmentor(1343)

Let M = the original amount
1st amount spent is 1/4M, so 3/4M remains
2nd amount spent is 1/5 of the remainder, or (1/5)(3/4M) = 3/20M.
The total amount spent is 1/4M + 3/20M = 8/20M = 2/5M.
3/5M = 66 -> M = $110